n Macintosh Ils a highly successful Pennsylvania orchardman who has formed his own company to produce and package Jonatha applesauce. Apples can be stored for several months in cold storage, so applesauce production is relotively uniform throughout the :. The recenthy hired controller for the firm ls about to apply the high-low method In estimating the companys energy cost behavior The following costs were incurred during the past 12 months: Pints of Applesauce January 116,000 74,000 77,000 83,000 101,000 107,000 131,000 101,000 $71,300 65,000 68,200 69,000 69,800 70,600 71,000 69,500 70,100 69,200 73,400 71,750 February March April May June July August September 101,000 October November 95,000 134,000 128,000 Required: 1. Use the high-low method to estimate the company bX, where Y denotes energy cost for a month and X denotes pints of applesauce produced. 2. Predict the energy cost for a month in which 85,000 pints of applesauce are produced. s energy cost behavior and express it in equation form. Use the formula Y at Complete this question by entering your answers in the tabs below Required 1 Required 2 Predict the energy cost for a month in which 85,000 pints of applesauce are produced S 66,540 K Prev 2 of 14E Next>
Answer #1

1.

Variable cost per unit = (Highest cost - Lowest cost)/Difference in output at highest and lowest cost

Variable energy cost = (73,400 - 65,000)/(134,000 - 74,000)

= 8,400/60,000

= 0.14 per pint

Hence, total variable cost for the month of February= 74,000 x 0.14

= $10,360

Fixed monthly cost = Total cost - Total variable cost

= 74,000 - 10,360

= $63,640

Cost function (Y) = a + bX

= 63,640 + 0.14X

2.

Energy cost of 85,000 pints = 63,640 + 0.14 x 85,000

= $75,540

Kindly give a positive rating if you are satisfied with the answer. Feel free to ask if you have any doubts. Thanks.